# Next Visit Diagnosis Prediction via Medical Code-Centric Multimodal Contrastive EHR Modelling with Hierarchical Regularisation

### v1

@@ A.3 Feature Extraction Module for Clinical Notes
First, we combine all notes $W_{t}^{1},W_{t}^{2},\ldots,W_{t}^{K}$ in a single patient visit $V_{t}$ to generate a single note $W_{t}$. Then, using the pre-trained BioWord2Vec (Zhang et al., 2019) $\text{E}_{\text{W}}$, each discrete word $w_{t}^{n}$ in the note $W_{t}$ is mapped to a low-dimensional embedding space, generating $e_{t}^{n}$. With the maximum number of words $|\mathbb{W}|$, the word embeddings $e_{t}=(e_{t}^{1},e_{t}^{2},\ldots,e_{t}^{|\mathbb{W}|})$ from the combined note $W_{t}$ are then fed into the 1D CNN (Conv1D) with multiple filters with a subsequent max-pooling layer (Max) to generate the most salient features $\bar{w}_{t}$ using a filter (equivalent to window size) $f$. The outputs from each filter are concatenated ($\oplus$) and passed to a linear layer with ReLU activation function. It yields the note representation $\bar{W}_{t}$ at $t$-th visit of each patient $\mathcal{P}$. The aforementioned processes are mathematically described as follows:
$W_{t}=W_{t}^{1}\oplus W_{t}^{2}\oplus,\cdots,\oplus W_{t}^{K},$ (26)
@@
$\bar{w}_{t}^{f}=\text{Max}(\bar{e}_{t}^{f})$ (29)
$\bar{w}_{t}=\bar{w}_{t}^{2}\oplus\bar{w}_{t}^{3}\oplus\bar{w}_{t}^{4},$ (30)
@@ Appendix B Data Pre-processing
We first pre-process the notes, following the previous work (Khadanga et al., 2019). It involves a removal of non-alphabetical characters, stopwords and conversion of uppercase to lowercase letters. Then, we add two special tokens to BioWord2Vec (Zhang et al., 2019), <UNK> and <PAD>, the same as those used in BERT (Devlin et al., 2018). They are initialised using matrices filled with zeros and uniform distribution, respectively. Any visit records lacking note information are excluded. Next, each note is tokenised with maximum 10k words using BioWord2Vec. This approach effectively captures the entirety of note information for approximately 85% of all the visits.
Medical Ontology & Label Construction Following the GRAM (Choi et al., 2017), a medical ontology is constructed based on ICD-9 codes using the Clinical Classifications Software (CCS) from the Healthcare Cost and Utilization Project [^2] https://hcup-us.ahrq.gov/toolssoftware/ccs/ccs.jsp. The labels are derived from nodes present in the primary and secondary hierarchy of the ICD-9 codes. This renders the next visit diagnosis prediction task as a hierarchical multi-label multi-class classification.
@@ D.1 Unimodal EHR Modelling Baselines
  • Demographics Encoder (Ours) utilises a simple combination of attribute-specific embedding layers and two linear layers, whose subsequent layers are ReLU and Sigmoid activation function, respectively. It is employed in our pipeline. For further details, refer to Appendix A.2.
  • BioWord2Vec (Zhang et al., 2019) model is combined with 1D CNN (Kim, 2014). For brevity, we simplify it as BioWord2Vec. It uses pre-trained BioWord2Vec embedding with 16,545,454 words (with an arbitrary addition of two special tokens), which are subsequently processed by 1D CNN. In our framework, this serves as the clinical notes feature extraction module. Refer to Appendix A.3 for details.
@@ D.2 Multimodal EHR Modelling Baselines
Both MNN and MAIN process 10k words from a clinical note within a single visit. The arameters (e.g. hidden dimension, the number of heads) are set in accordance with the specifications detailed in their original paper.
@@ Appendix E A Comparative Study on Different MAGs
From the Table 6, it demonstrates the superiority of our method over preceding approaches. It can be attributed to the meticulous consideration of the modality imbalance, one of factors not adequately addressed by previous methodologies. This validates that considering dominance of main modality is essential in multimodal modelling.

### v5

@@ A.3 Feature Extraction Module for Clinical Notes
First, we combine all notes $W_{t}^{1},W_{t}^{2},\ldots,W_{t}^{K}$ in a single patient visit $V_{t}$ to generate a single note $W_{t}$. Then, using the pre-trained BioWord2Vec (Zhang et al., 2019) $\text{E}_{\text{W}}$, each discrete word $w_{t}^{n}$ in the note $W_{t}$ is mapped to a low-dimensional embedding space, generating $e_{t}^{n}$. With the maximum number of words $|\mathbb{W}|$, the word embeddings $e_{t}=(e_{t}^{1},e_{t}^{2},\ldots,e_{t}^{|\mathbb{W}|})$ from the combined note $W_{t}$ are then fed into the 1D CNN (Conv1D) with multiple filters with a subsequent max-pooling layer (Max) to generate the most salient features $\bar{w}_{t}$ using a filter (equivalent to window size) $f$. The outputs from each filter are concatenated and passed to a linear layer with ReLU activation function. It yields the note representation $\bar{W}_{t}$ at $t$-th visit of each patient $\mathcal{P}$. The aforementioned processes are mathematically described as follows:
$W_{t}=\text{concat}(W_{t}^{1};W_{t}^{2};\cdots;W_{t}^{K}),$ (26)
@@
$\bar{w}_{t}^{f}=\text{Max}(\bar{e}_{t}^{f}),$ (29)
$\bar{w}_{t}=\text{concat}(\bar{w}_{t}^{2};\bar{w}_{t}^{3};\bar{w}_{t}^{4}),$ (30)
@@ Appendix B Data Pre-processing
We first pre-process the notes, following the previous work (Khadanga et al., 2019). It involves a removal of non-alphabetical characters, stopwords and conversion of uppercase to lowercase letters. Then, we add two special tokens to BioWord2Vec (Zhang et al., 2019), <PAD> and <UNK>, the same as those used in BERT (Devlin et al., 2018). They are initialised using matrices filled with zeros and uniform distribution, respectively. Any visit records lacking note information are excluded. Next, each note is tokenised with maximum 10k words using BioWord2Vec. This approach effectively captures the entirety of note information for approximately 85% of all the visits.
Medical Ontology & Label Construction Following the GRAM (Choi et al., 2017), a medical ontology is constructed based on ICD-9 codes using the Clinical Classifications Software (CCS) from the Healthcare Cost and Utilization Project [^1] https://hcup-us.ahrq.gov/toolssoftware/ccs/ccs.jsp. The labels are derived from nodes present in the primary [^2] https://hcup-us.ahrq.gov/toolssoftware/ccs/AppendixCMultiDX.txt and secondary [^3] https://hcup-us.ahrq.gov/toolssoftware/ccs/AppendixASingleDX.txt hierarchy of the ICD-9 codes. This renders the next visit diagnosis prediction task as a hierarchical multi-label multi-class classification.
@@ D.1 Unimodal EHR Modelling Baselines
  • Medical Code Encoder (Ours) employs a simple combination of embedding layers and a couple of linear layers, which are followed by ReLU and Sigmoid activation function. It is utilised in our pipeline. Refer to Appendix A.1 for details.
  • Demographics Encoder (Ours) utilises a simple combination of attribute-specific embedding layers and two linear layers, whose subsequent layers are ReLU and Sigmoid activation function, respectively. It is employed in our pipeline. Refer to Appendix A.2 for details.
  • BioWord2Vec (Zhang et al., 2019) model is combined with 1D CNN (Kim, 2014). For brevity, we simplify it as BioWord2Vec. It uses pre-trained embedding with 16,545,454 words (with an arbitrary addition of two special tokens), which are subsequently processed by 1D CNN. In our framework, this serves as the notes feature extraction module. Refer to Appendix A.3 for details.
@@ D.2 Multimodal EHR Modelling Baselines
Both MNN and MAIN process 10k words from a clinical note within a single visit. The parameters (e.g. hidden dimension, the number of heads) are set in accordance with the specifications detailed in their original paper.
@@ Appendix E A Comparative Study on Different MAGs
From the Table 6, it demonstrates the superiority of our method over preceding approaches. It can be attributed to the meticulous consideration of the modality imbalance, one of factors not adequately addressed by previous methodologies. This validates that considering the dominance of main modality is essential in multimodal modelling.